3x + 4y = 12 (0, ),( ,3/4),( ,0),(8/3, )
Answer by funmath(2933)

3x + 4y = 12 (0, ),( ,3/4),( ,0),(8/3, )
Ordered pairs are in this order (x,y)
For (0, ), let x=0 and solve for y.
3(0)+4y=12
0+4y=12
4y=12
4y/4=12/4
y=3
Therefore, (x,y)=(0,3)
:
For ( ,3/4), let y=3/4 and solve for x.
3x+4(3/4)=12
3x+12/4=12
3x+3=12
3x+3-3=12-3
3x=9
3x/3=9/3
x=3
Therefore, (x,y)=(3,3/4)
:
For ( ,0), let y=0 and solve for x.
3x+4(0)=12
3x+0=12
3x=12
3x/3=12/3
x=4
Therefore, (x,y)=(4,0)
:
For (8/3, ), let x=8/3 and solve for y.
3(8/3)+4y=12
24/3+4y=12
8+4y=12
8-8+4y=12-8
4y=4
4y/4=4/4
y=1
Therefore, (x,y)=(8/3,1)
Happy Calculating!!!